twice the difference of a number and 8 is equal to three times the sum of the numer and 3 find the number?
x-2+8 not sure if thats correct but my text book got -25 as an answer can anyone help me with how my text book got -25

Found 2 solutions by mananth, haileytucki:
Answer by mananth(16946)

let number be x
twice the difference between the number and 8 = 2(x-8)
3 times sum of number and 3 = 3(x+3)
2(x-8)=3(x+3)
2x-16=3x+9
rearrange the terms and numbers
3x-2x= -16-9
x= -25

Answer by haileytucki(390)

I believe this is what you are needing:
2x-x+8=3x+3
Since 2x and -x are like terms, add -x to 2x to get x.
x+8=3x+3
Since 3x contains the variable to solve for, move it to the left-hand side of the equation by subtracting 3x from both sides.
x+8-3x=3
Since x and -3x are like terms, add -3x to x to get -2x.
-2x+8=3
Since 8 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 8 from both sides.
-2x=-8+3
Add 3 to -8 to get -5.
-2x=-5
Divide each term in the equation by -2.
-(2x)/(-2)=-(5)/(-2)
Simplify the left-hand side of the equation by canceling the common factors.
x=-(5)/(-2)
x= -25